OpenStudy (anonymous):

Explain how to express series using sigma notation : 4+8+12+16+20+...

OpenStudy (anonymous):

\[\sum_{k=1}^{\infty} \] what I have so far ^.

OpenStudy (anonymous):

\[\sum_{n=1}^{n=\infty} 4n\]

OpenStudy (anonymous):

how did you figure that out ?

OpenStudy (anonymous):

see all the terma are multiples of 4 4*1=4 4*2=8 and so on......... summation means to sum up........so. for n=1,we get 4 .......then for n=2,we get 8.......for n=3,we get 12.....and so on so ,final result 4+8+12+16+20+...

OpenStudy (anonymous):

its terms........

OpenStudy (anonymous):

didnt get it???

OpenStudy (anonymous):

got it. thanks